if f(x)=x^2+5, find f(a+h)-f(a)
Answer by Nate(3500)

f(x) = x^2 + 5
f(a + h) = (a + h)^2 + 5 = a^2 + 2ah + h^2 + 5
f(a) = (a)^2 + 5 = a^2 + 5
~~
f(a + h) - f(a)
(a^2 + 2ah + h^2 + 5) - (a^2 + 5)
a^2 + 2ah + h^2 + 5 - a^2 - 5
2ah + h^2